# Spatiotemporal Photon Blockade for Nonreciprocal Quantum Absorption

### v3

@@ 1 Spatiotemporal Photon Blockade
Figure 2 illustrates the proposed one-way photon blockade and absorption mechanism facilitated by the coherently driven quantum spatiotemporal metasurface and its functionality. The metasurface operates as a nonreciprocal element, permitting photon transmission in one direction while blocking it in the opposite direction. To understand the nonreciprocal behavior of the spatiotemporal quantum metasurface, we first analyze its band structure using the dispersion relation. The analysis reveals crucial insights into the energy transitions and nonreciprocity induced by the space-time modulation. Figure 3a presents the $\omega-\kappa$ diagram, showcasing the band structure of the metasurface at the critical point where $\omega_{0}/\omega_{\text{s}}=1$. This diagram illustrates a strong nonreciprocity in the system. Specifically, the higher-order harmonics that travel in the forward $+z$ direction converge at the normalized wave vector $\kappa_{n}/\kappa_{\text{s}}=1$. This convergence indicates a resonant interaction between the incident photons and the space-time-modulated medium, leading to the absorption of energy from the fundamental harmonic. The absorbed energy causes a transition from the fundamental frequency $\omega_{0}$ to higher energy states denoted as $\omega_{m}$.
@@
Figure 3: Band structure of the spatiotemporal quantum metasurface computed. (a) The $\omega-\kappa$ diagram illustrates strong nonreciprocity at $\omega_{0}/\omega_{\text{s}}=1$. Here, all higher-order forward ($+z$ direction) traveling space-time harmonics converge at $\kappa_{n}/\kappa_{\text{s}}=1$, resulting in the absorption of energy from the fundamental harmonic and transitioning the system from $\omega_{0}$ to higher frequency states, denoted as $\omega_{m}$. In contrast, the higher-order backward ($-z$ direction) traveling space-time harmonics are spatially separated, preventing any energy transition from $\omega_{0}$ to higher states. (b) The $k_{x}-\kappa_{n}$ isofrequency diagram at $\omega_{0}/\omega_{\text{s}}=1$ further demonstrates the nonreciprocity of the metasurface, where all forward higher-order harmonics are clustered at $\kappa_{n}/\kappa_{\text{s}}=1$, effectively absorbing the energy of the fundamental harmonic. The group velocity vector $v_{\text{g},n}$ indicates that these harmonics propagate along the $+z$ direction, parallel to the metasurface boundary, rather than being transmitted outside the metasurface.
@@
In contrast, the higher-order harmonics that travel in the backward $-z$ direction are spatially separated in the $\omega-\kappa$ diagram. This separation prevents any significant interaction between the incident photons and the medium, thereby inhibiting the energy transition from $\omega_{0}$ to higher states. As a result, the backward-propagating waves do not experience the same degree of energy absorption as the forward-propagating waves, which directly contributes to the observed nonreciprocal behavior of the metasurface. Figure 3b shows the $k_{x}-\kappa_{n}$ isofrequency diagram at $\omega_{0}/\omega_{\text{s}}=1$. This diagram further confirms the nonreciprocal nature of the metasurface, where all forward-propagating higher-order harmonics are clustered at $\kappa_{n}/\kappa_{\text{s}}=1$. The clustering of these harmonics facilitates the absorption of the fundamental harmonic’s energy, while the group velocity vector $v_{\text{g},n}$ indicates that the energy is predominantly confined and directed along the $+z$ axis, parallel to the metasurface boundary. This confinement prevents the transmission of energy to the exterior of the metasurface, leading to strong nonreciprocal absorption.
@@
The phenomena observed in the band structure is related to the concept of spatiotemporal photon blockade. In photon blockade, the presence of a single photon can prevent the passage of subsequent photons due to nonlinear interactions, effectively ”blocking” further photon transmission. In our system, this blockade effect is realized through the spatiotemporal modulation of the metasurface. The space-time modulation, with a frequency $\omega_{\text{s}}$ matched to the incident photon frequency $\omega_{0}$, induces a strong nonlinear interaction for forward-propagating waves. As a result, when photons from the left are incident on the metasurface, they are resonantly coupled to the modulation, leading to energy absorption and transition to higher frequency states (spatial harmonics) $\omega_{m}$. This absorption effectively prevents the transmission of these photons through the metasurface, analogous to a photon blockade. The nonreciprocity arises because the same modulation does not facilitate similar energy transitions for backward-propagating waves, allowing them to pass through the metasurface without significant energy loss. Thus, the metasurface behaves as a nonreciprocal photon blockade device, where the space-time modulation selectively blocks photons traveling in one direction by shifting their energy to higher states, while allowing photons traveling in the opposite direction to pass freely. This behavior is crucial for potential applications in quantum information processing and nonreciprocal quantum devices, where controlling the directionality and energy states of photons is essential.
Next, we analyze the nonreciprocal absorption characteristics of the spatiotemporal quantum metasurface. The metasurface is designed to exhibit strong nonreciprocal behavior due to the space-time modulation, leading to direction-dependent absorption and transmission properties. Figures 4a and 4b illustrate the electric field distribution ($E_{y}$) and corresponding frequency spectrum for two different incident angles. Specifically, Figure 4a demonstrates the field distribution and frequency response when an electromagnetic wave is incident from the left at an angle of $55^{\circ}$. The results clearly show strong absorption of the incident wave by the metasurface, with minimal transmission, indicating effective photon blockade in this direction. The corresponding frequency spectrum confirms the absorption by showing the lack of transmitted energy at the incident frequency. Conversely, Figure 4b presents the scenario where the incident beam approaches from the right at an angle of $125^{\circ}$. In this case, the metasurface allows full transmission of the wave to the left side, effectively reflecting the wave back towards the source at an angle of $55^{\circ}$. The frequency spectrum corroborates this by displaying a significant transmitted signal at the original frequency, confirming the nonreciprocal behavior of the metasurface. Figures 4c and 4d provide the magnetic field distribution ($H_{z}$) under the same conditions. Figure 4c shows the absorption of the left-incident wave at $55^{\circ}$, while Figure 4d demonstrates the full transmission of the right-incident wave at $125^{\circ}$. The consistent behavior observed in both the electric and magnetic field distributions confirms the robustness of the nonreciprocal absorption mechanism induced by the spatiotemporal modulation in the superconducting quantum metasurface. These results demonstrate the capability of the designed metasurface to enforce nonreciprocal photon blockade, effectively controlling the directionality of absorption and transmission. The strong interaction between the incident photons and the spatiotemporally modulated nonlinear metasurface unit cells underpins the observed nonreciprocal behavior, which could be leveraged in various millikelvin-temperature quantum technology applications requiring directional control of photon propagation.
@@
Figure 4: Nonreciprocal absorption in spatiotemporal quantum metasurfaces. (a) and (b) Field distribution (top) and frequency spectrum (bottom) for $E_{y}$, demonstrating (a) strong absorption of the incident beam from the left at $55^{\circ}$, and (b) full transmission of the incident beam from the right at $125^{\circ}$ to the left at $55^{\circ}$. (c) and (d) Field distribution for $H_{z}$, demonstrating (c) strong absorption of the incident beam from the left at $55^{\circ}$, and (d) full transmission of the incident beam from the right at $125^{\circ}$ to the left at $55^{\circ}$.

### v4

@@ 3.2 Dispersion analysis
To understand the nonreciprocal behavior of the superconductor-semiconductor spatiotemporal quantum metasurface, we first analyze its band structure using the dispersion relation. The analysis reveals crucial insights into the energy transitions and nonreciprocity induced by the space-time modulation. Figure 4a presents the $\omega-\kappa$ diagram, showcasing the band structure of the metasurface at the critical point where $\omega_{0}/\omega_{\text{s}}=1$. This diagram illustrates a strong nonreciprocity in the system. Specifically, the higher-order harmonics that travel in the forward +z+z direction converge at the normalized wave vector $\kappa_{n}/\kappa_{\text{s}}=1$. This convergence indicates a resonant interaction between the incident wave and the space-time-modulated medium, leading to the absorption of energy from the fundamental harmonic. The absorbed energy causes a transition from the fundamental frequency $\omega_{0}$ to higher energy states denoted as $\omega_{m}$.
@@
Figure 4: Band structure of the superconductor-semiconductor spatiotemporal quantum metasurface computed using Eq. (22). (a) The $\omega-\kappa$ diagram illustrates strong nonreciprocity at $\omega_{0}/\omega_{\text{s}}=1$. Here, all higher-order forward (+z+z direction) traveling space-time harmonics converge at $\kappa_{n}/\kappa_{\text{s}}=1$, resulting in the absorption of energy from the fundamental harmonic and transitioning the system from $\omega_{0}$ to higher frequency states, denoted as $\omega_{m}$. In contrast, the higher-order backward (z-z direction) traveling space-time harmonics are spatially separated, preventing any energy transition from $\omega_{0}$ to higher states. (b) The $k_{x}-\kappa_{n}$ isofrequency diagram at $\omega_{0}/\omega_{\text{s}}=1$ further demonstrates the nonreciprocity of the metasurface, where all forward higher-order harmonics are clustered at $\kappa_{n}/\kappa_{\text{s}}=1$, effectively absorbing the energy of the fundamental harmonic. The group velocity vector $v_{\text{g},n}$ indicates that these harmonics propagate along the +z+z direction, parallel to the metasurface boundary, rather than being transmitted outside the metasurface.
@@
In contrast, the higher-order harmonics that travel in the backward z-z direction are spatially separated in the $\omega-\kappa$ diagram. This separation prevents any significant interaction between the incident wave and the medium, thereby inhibiting the energy transition from $\omega_{0}$ to higher states. As a result, the backward-propagating waves do not experience the same degree of energy absorption as the forward-propagating waves, which directly contributes to the observed nonreciprocal behavior of the metasurface. Figure 4b shows the $k_{x}-\kappa_{n}$ isofrequency diagram at $\omega_{0}/\omega_{\text{s}}=1$. This diagram further confirms the nonreciprocal nature of the metasurface, where all forward-propagating higher-order harmonics are clustered at $\kappa_{n}/\kappa_{\text{s}}=1$. The clustering of these harmonics facilitates the absorption of the fundamental harmonic’s energy, while the group velocity vector $v_{\text{g},n}$ indicates that the energy is predominantly confined and directed along the +z+z axis, parallel to the metasurface boundary. This confinement prevents the transmission of energy to the exterior of the metasurface, leading to strong nonreciprocal absorption.
@@
The phenomena observed in the band structure is related to the concept of spatiotemporal photon blockade, where the presence of a single photon can prevent the passage of subsequent photons due to nonlinear interactions, effectively ”blocking” further photon transmission. In our system, this effect is realized through the spatiotemporal modulation of the metasurface. The space-time modulation, with a frequency $\omega_{\text{s}}$ matched to the incident photon frequency $\omega_{0}$, induces a strong nonlinear interaction for forward-propagating waves. As a result, when photons from the left are incident on the metasurface, they are resonantly coupled to the modulation, leading to energy absorption and transition to higher frequency states (spatial harmonics) $\omega_{m}$. This absorption effectively prevents the transmission through the metasurface. The nonreciprocity arises because the same modulation does not facilitate similar energy transitions for backward-propagating waves, allowing them to pass through the metasurface without significant energy loss. Thus, the metasurface behaves as a nonreciprocal device, where the space-time modulation selectively blocks traveling in one direction by shifting their energy to higher states, while allowing wave traveling in the opposite direction to pass freely. This behavior is crucial for potential applications in quantum information processing and nonreciprocal quantum devices, where controlling the directionality and energy states of photons is essential.
3.3 Transmission and fields distributions
@@
Figure 5: Nonreciprocal absorption in spatiotemporal superconductor-semiconductor quantum metasurfaces. (a) and (b) Field distribution (top) and frequency spectrum (bottom) for $E_{y}$, demonstrating (a) strong absorption of the incident beam from the left at $55^{\circ}$, and (b) full transmission of the incident beam from the right at $125^{\circ}$ to the left at $55^{\circ}$. (c) and (d) Field distribution for $H_{z}$, demonstrating (c) strong absorption of the incident beam from the left at $55^{\circ}$, and (d) full transmission of the incident beam from the right at $125^{\circ}$ to the left at $55^{\circ}$.
Next, we analyze the nonreciprocal absorption characteristics of the spatiotemporal superconductor-semiconductor quantum metasurface. The metasurface is designed to exhibit strong nonreciprocal behavior due to the space-time modulation, leading to direction-dependent absorption and transmission properties. Figures 5a and 5b illustrate the electric field distribution ($E_{y}$) and corresponding frequency spectrum for two different incident angles. Specifically, Figure 5a demonstrates the field distribution and frequency response when an electromagnetic wave is incident from the left at an angle of $55^{\circ}$. The results clearly show strong absorption of the incident wave by the metasurface, with minimal transmission. The corresponding frequency spectrum confirms the absorption by showing the lack of transmitted energy at the incident frequency. Conversely, Figure 5b presents the scenario where the incident beam approaches from the right at an angle of $125^{\circ}$. In this case, the metasurface allows full transmission of the wave to the left side, effectively reflecting the wave back towards the source at an angle of $55^{\circ}$. The frequency spectrum corroborates this by displaying a significant transmitted signal at the original frequency, confirming the nonreciprocal behavior of the metasurface. Figures 5c and 5d provide the magnetic field distribution ($H_{z}$) under the same conditions. Figure 5c shows the absorption of the left-incident wave at $55^{\circ}$, while Figure 5d demonstrates the full transmission of the right-incident wave at $125^{\circ}$. The consistent behavior observed in both the electric and magnetic field distributions confirms the robustness of the nonreciprocal absorption mechanism induced by the spatiotemporal modulation in the superconducting quantum metasurface. These results demonstrate the capability of the designed metasurface to enforce nonreciprocity, effectively controlling the directionality of absorption and transmission. The strong interaction between the incident photons and the spatiotemporally modulated nonlinear metasurface unit cells underpins the observed nonreciprocal behavior, which could be leveraged in various millikelvin-temperature quantum technology applications requiring directional control of wave propagation.